OpenStudy (anonymous):

How do you rewrite an equation from point-slope form to slope-intercept form? And from slope-intercept form, how do you convert to standard form?

OpenStudy (anonymous):

If I remember it properly point-slope for is in terms of x= and slope-intercept form is in terms of y=

hero (hero):

Begin by writing an equation in point slope form: (y - y1) = m(x - x1). Then manipulate it until you have the form y = mx + b. Then, go a step further and manipulate it until you have the form ax + by = c For example, suppose you're given the point (x,y) = (2,5) and the slope m = 3. Rewriting this in point slope form produces: (y - 2) = 3( x -5) . To re-write this equation in slope-intercept form, manipulate the equation as follows: (y - 2) = 3(x - 5) y - 2 = 3x - 15 y = 3x - 15 + 2 y = 3x - 13 This last form, y = 3x - 13 is the slope-intercept form y = mx + b. Finally, to put this equation in standard form ax + by = c, manipulate the equation as follows: y = 3x - 13 13 = 3x - y 13 = 3x - 1y 3x + (-1)y = 13 The last form, 3x + (-1)y = 13 is the standard form ax + by = c.